OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y ≤ x - 5 y ≥ -x - 4 a) (-5, 2) b) (5, -2) c) (-5, -2) d) (5, 2)

OpenStudy (anonymous):

a) and c) are ruled out because if x = -5.. y <= x - 5 y <= (-5) - 5 y <= -10 but the y coordinate in a) is 2, and in c.) is -2. both of which are greater than -10 d is ruled out for the same reason if x = 5 y <= x - 5 y <= 5 - 5 y <= 0 but d.) has a y component of 2, which is greater than 0 that leaves b.)